OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. -5/7-2/5

OpenStudy (anonymous):

\[\LARGE -\frac57 -\frac25=-\frac{5\cdot 5}{7\cdot 5}-\frac{2\cdot ?}{5\cdot 7}=? \]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

"that's not the answer " .. is juts a hint... can you continue ?

OpenStudy (anonymous):

yes I have it from here!! thanks :)

OpenStudy (anonymous):

wait... is thetop number 5? If not... I am way off track

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

it's 5 and that becomes.. \[\LARGE -\frac{25}{35}-\frac{10}{35}=-\frac{(25+10)}{35}=?\]